# Ziniu Hu4 , Haifeng Chen3,  Wei Cheng🖂3 1

### v1

@@ 1 Introduction
Specifically, SFS contains three key techniques. Scattering is a novel technique that dynamically varies input prompts when sampling from LLMs, driving more diverse and exploratory outputs. In Scattering, the LLM suggests different textual optimization directions and steps, analogous to gradients in numerical optimization, before advancing towards a new solution. During tree search refinement, Scattering effectively perturbs or mutates previous solutions, resulting in an evolutionary search process. We further propose Foresting, the tree search equivalent of multi-start optimization, where Scattering plays a crucial role in diversifying initialization seeds, ensuring they are well-distributed throughout the search space. This enhances the breadth of exploration while effectively mitigating clearly incorrect solutions, such as those containing syntax errors.
@@
    We analyze the lack of diversity in current LLM code search processes and frame code generation as a black-box optimization problem within the language space, emphasizing the need for a balance between exploration and exploitation.
@@
    We demonstrate that our method significantly enhances the accuracy, scalability, and solution diversity of the search process across widely used code generation benchmarks, including HumanEval, MBPP, Leetcode, APPS, and CodeContests.
@@ 2.1 Problem description
In the task $\bm{x}=\langle\bm{p},\bm{H}\rangle$ of program synthesis, the solver is given a prompt $\bm{p}$ in natural language, which asks the solver to write code for some object $\bm{s}$. The goal is to complete the code implementation of $\bm{s}$ such that it passes all the hidden tests $\bm{H}$. The solver is not allowed to see the hidden tests. Sometimes, the solver is also given validation (visible) tests $\bm{V}$ that they can test their solution $\bm{s}$ on before they submit it for evaluation on the hidden tests $\bm{H}$. They can also generate their own validation tests $\bm{V}$. Usually both hidden and validation tests are in the form of a set of assert statements that test the functionalities of $\bm{s}$. An example of one such code generation task is shown in Appendix B.
@@ 3 Methodology
Our method incorporates three optimization inspired techniques to enhance both exploration and exploitation of tree search (MCTS) using LLMs.
Figure 7: Core techniques used by SFS. Points represent solutions. Hexagons represent initial solutions. Star represents the final selected solution.
@@ 3.1 Tree branch Scattering
$UCT(\bm{s},\bm{d})=\widehat{Q}(\bm{s},\bm{d})+c\sqrt{\frac{\ln\left(\sum_{\bm{b}}n(\bm{s},\bm{d})\right)}{n(\bm{s},\bm{d})}},$ (1)
@@
$\widehat{Q}(\bm{s}_{i},\bm{d}_{i})^{(t+1)}\leftarrow(1-\alpha(n))\widehat{Q}(\bm{s}_{i},\bm{d}_{i})^{(t)}+\alpha(n)\mathop{max}\{\widehat{Q}(\bm{s}_{i},\bm{d}_{i})^{(t)},\widehat{Q}(\bm{s}_{ij},\bm{d}_{ij})^{(t+1)}\},$ (2)
where $\alpha(n)$ is the weighted average parameter that depends on $n(\bm{s},\bm{d})$. The backpropogation occurs along the entire MCTS simulated trajectory $\tau=[\bm{s}_{0},\bm{d}_{0i},\bm{s}_{0i},...,\bm{s}_{-2},\bm{d}_{-1},\bm{s}_{-1}]$, where the $q$-value for the penultimate state $\bm{s}_{-1}$ is updated using the evaluation value of the final state $\widehat{Q}(\bm{s}_{-2},\bm{d}_{-1})^{(t+1)}\leftarrow v(\bm{s}_{-1})$ instead. We take the max of $\widehat{Q}(\bm{s}_{i},\bm{d}_{i})^{(t)}$ and the next state’s $q$-value $\widehat{Q}(\bm{s}_{ij},\bm{d}_{ij})$ to ensure that if the next solution is worse, the current solution can be used instead. This approach dynamically selects which direction to explore, prioritizing more promising parent solutions $\bm{s}_{j}$ over exploring all directions of a parent solution $\bm{s}_{i}$. Using UCT to select distinct actions has been effective in balancing exploration and exploitation in other settings (Browne et al., 2012).
@@ 3.3 Branch Scouting
Inspired by traditional optimization techniques like Ant Colony Optimization and Particle Swarm optimization, we utilize insights from other branches—specifically, which improvement directions are effective—to enhance the current solution. Specifically, after generating a new solution $\bm{d}_{-1}$ by applying improvement $\bm{d}_{-1}$ to solution $\bm{d}_{-2}$, we provide feedback on the new solution $\bm{f}_{-1}$ to the LLM, inquiring whether the improvement was effective and what general insights it can derive from this. The insights will then be store in global memory, and in future iterations when the LLM is prompted to generate improvement directions, the global insights will be included in the prompt. This approach enables knowledge about effective and ineffective improvement directions to be shared across branches, facilitating more efficient exploitation of the search feedback and thereby enhancing our Scattering technique.
@@ 3.4 A Theoretical Perspective
In Markov chain theory, a diverse transition kernel increases the probability of moving between different regions of the state space (denoted by $\mathcal{S}$), thereby enhancing the conductance $\Phi$ of the chain:
@@
where $S$ is a subset and $\mu$ is the stationary distribution.
@@ 3.5 Empirical validation
In addition to the metrics we discussed previously, we also include other similarity metrics such as tf-idf similarity, which measures the average cosine similarity between tf-idf vectors, the Levenshtein similarity, and the token sequence similarity. See App. I for more details, including examples.
@@ 4.2 Accuracy
LATS111ran under our setup, see App. D for similar setup  (Zhou et al., 2024) $75.6$ $79.6$
Ours (SFS) 82.5 81.7
@@ 4.3 Scalability
Figure 12: Scaling curves for different search methods on CodeContests (left) and Leetcode (right). We run each method for 10 iterations using gpt-3.5, reporting the proportion of problems where the correct solution is discovered at each iteration. Additional curves can be found in Sec. C
We report the average number of iterations (solutions generated) it takes before the search algorithm discovers the correct solution in Table 6. Iters. (incl) is the average number of iterations it takes including problems where the algorithm succeeds on the first try. Iters. (excl) is the average number of iterations it takes excluding first try successes, where the search algorithm is actually used to find the correct solution. On both metrics, our method demonstrates the ability to discover the correct solution much more quickly than the other methods. Moreover, we see in Figure 2 and 12 that our method also scales better than the other methods on all datasets.
Table 6: Metrics for different search methods. We run different search methods for 10 iterations each using gpt-3.5-turbo on HumanEval. Our method generates more diverse solutions and discovers the correct solution faster.
@@ 4.4 Solution diversity
We see in Table 6 that our proposed search method is able to propose more diverse candidate solutions with a lower semantic similarity score, while maintaining high quality search with by generating solutions with high validation scores. This shows that our methods help both exploration and exploitation without sacrificing too much of one for the other. Detailed stats shown in App. F.
@@ 4.5 Ablation on techniques
We performed an ablation study on the three introduced techniques as shown in Table 7 (and App. G), all of which enhanced performance and efficiency, with Scattering yielding the highest gains.
@@ 4.6 Verifier (Validation test) accuracy
Previous work emphasized the importance of a reliable verifier  (Liu et al., 2024a; Chen et al., 2022; Zhang et al., 2023a). In our case, we use self-generated validation tests noisy as a black-box verifer. Figure 13 shows the confusion matrix for whether self-generated validation tests accurately predict solution correctness. The false negative rate of 27.5% highlights misalignment between validation and ground truth tests. We performed an ablation study comparing performance when given different numbers of ground-truth tests as validation (Table 8). While finding the best verifier is not our focus, the results show a significant impact of verifier accuracy on performance. There are strong interactions between the search method and evaluation metric—just one ground-truth validation test dramatically improves performance. This demonstrates how well our methods exploit accurate evaluation signals and perform even better when the black-box evaluation is more accurate. Detailed metrics shown in App. H.
@@ 4.7 Ablation on model
As shown in Figure 14, weaker models scale better with our method. This supports the trade-off between training and inference compute: better-trained models scale worse with inference compute, and vice versa. Detailed stats in App. E.
@@ 5 Additional Related work
Code generation with large language models. Some works focus on either training their own model  (Guo et al., 2024) or fine tuning existing models to adapt them towards code related tasks  (Roziere et al., 2023; Jain et al., 2023; Roziere et al., 2023). Recent literature has shown a flora of methods to improve LLM code generation performance during inference time, including agentic approaches  (Qian et al., 2023; Liu et al., 2024b), multi-agent approaches  (Tao et al., 2024; Hong et al., 2023; Islam et al., 2024), tool usage  (Zhang et al., 2024), and self-revision  (Le et al., 2023). Our work compliments these works by showing that increasing solution diversity and exploration can have a large impact on inference scaling.
@@ 6 Conclusion
[…]
The integration of search-based methods could significantly reduce computational costs while maintaining or enhancing performance, making them attractive for large-scale deployments, especially in real-time applications or those requiring resource efficiency. This framework also lays the groundwork for future research into optimization strategies for language models, potentially leading to more advanced search algorithms, hybrid models, and novel techniques that push the limits of generative models.
@@ Appendix A Ethics Statement
Appendix B Coding problem example
@@
Appendix C Scaling
Our method scales well with increased iterations. We show scaling curves for each dataset as shown in Figures 15, 16, 17.

### v2

@@ 1 Introduction
[…]
@@
    We frame code generation as a black-box optimization problem over the code/text space and demonstrate how gradient-free optimization techniques can be adapted to enhance inference scaling. This perspective highlights the importance of balancing exploration and exploitation in search-based code generation and the need for diverse generation.
@@
    We empirically validate our approach across multiple code generation benchmarks, including HumanEval, MBPP, Leetcode, APPS, and CodeContests, demonstrating significant improvements in accuracy, scalability, and solution diversity over existing search methods.
@@ 2.1 Problem description
In a program synthesis task $\bm{x}=\langle\bm{p},\bm{H}\rangle$, the solver is given a prompt $\bm{p}$ in natural language, which asks the solver to write code for some object $\bm{s}$. The goal is to complete the code implementation of $\bm{s}$ such that it passes all the hidden tests $\bm{H}$. The solver is not allowed to see the hidden tests. Sometimes, the solver is also given validation (visible) tests $\bm{V}$ that they can test their solution $\bm{s}$ on before they submit it for evaluation on the hidden tests $\bm{H}$. They can also generate their own validation tests $\bm{V}$. Usually both hidden and validation tests are in the form of a set of assert statements that test the functionalities of $\bm{s}$. A description of one such code generation task is shown in Appendix B.
@@ 3 Methodology
Our method incorporates three optimization inspired techniques to enhance both exploration and exploitation of tree search (MCTS) using LLMs. More details in Appendix C with pseudocode.
[…]
@@ 3.1 Tree branch Scattering
$UCT(\bm{s},\bm{d})=\widehat{Q}(\bm{s},\bm{d})+c\sqrt{\frac{\ln\left(\sum_{\bm{b}}n(\bm{s},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{b}})\right)}{n(\bm{s},\bm{d})}},$ (1)
@@
$\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n}\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)}+\alpha_{n}\mathop{max}\{\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)},\widehat{Q}(\bm{s}_{i+1},\bm{d}_{i+2})^{(t+1)}\}$ (2)
where $\alpha_{n}$ is the weighted average parameter that depends on $n(\bm{s},\bm{d})$. The backpropogation occurs along the entire MCTS simulated trajectory $\tau^{(t+1)}=[\bm{s}_{0},\bm{d}_{1},\bm{s}_{2},...,\bm{s}_{-2},\bm{d}_{-1},\bm{s}_{-1}]$, where the $q$-value for the penultimate state $\bm{s}_{-1}$ is updated using the value of the final state (of validation tests passed) $\widehat{Q}(\bm{s}_{-2},\bm{d}_{-1})^{(t+1)}\leftarrow v(\bm{s}_{-1})$. We take the max of $\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)}$ and $\widehat{Q}(\bm{s}_{i+1},\bm{d}_{i+2})^{(t+1)}$ to ensure that if the next solution is worse, the current solution can be used instead. This approach dynamically selects which direction to explore, prioritizing more promising parent solutions $\bm{s}_{j}$ over exploring all directions of a parent solution $\bm{s}_{i}$. Using UCT to select distinct actions has been effective in balancing exploration and exploitation in other settings (Browne et al., 2012).
@@ 3.3 Branch Scouting
Inspired by optimization techniques like Ant Colony Optimization and Particle Swarm Optimization, we improve solutions by leveraging insights from effective improvement directions. After generating a new solution $\bm{s}_{-1}$ using improvement directions $\bm{d}_{-1}$ on $\bm{s}_{-2}$, we provide feedback $\bm{f}_{-1}$ to the LLM to assess its effectiveness and derive general insights. These insights are stored in global memory and included in future prompts, enabling shared knowledge of effective improvements across branches. This enhances feedback exploitation and strengthens our Scattering technique.
@@ 3.4 A Theoretical Perspective
In Markov chain theory, a diverse transition kernel increases the probability of moving between different regions of the state space (denoted by $\mathcal{S}$), enhancing the conductance $\Phi$ of the chain:
@@
where $S$ is a subset and $\mu$ is the stationary distribution. More details in Appendix M.
@@ 3.5 Empirical validation
In addition to the metrics we discussed previously, we also include other similarity metrics such as tf-idf similarity, which measures the average cosine similarity between tf-idf vectors, the Levenshtein similarity, and the token sequence similarity. See App. J for more details, including examples.
@@ 4.2 Accuracy
LATS111ran under our setup, see App. E for similar setup  (Zhou et al., 2024) $75.6$ $79.6$
[…]
@@ 4.3 Scalability
[…]
We report the average number of iterations (solutions generated) it takes before the search algorithm discovers the correct solution in Table 6. Iters. (incl) is the average number of iterations it takes including problems where the algorithm succeeds on the first try. Iters. (excl) is the average number of iterations it takes excluding first try successes, where the search algorithm is actually used to find the correct solution. On both metrics, our method demonstrates the ability to discover the correct solution much more quickly than the other methods. Moreover, we see in Figure 2 and 12 that our method also scales better than the other methods on all datasets. Our method scales efficiently with higher budgets, improving up to 20 solutions while others plateau as shown in Figure 18.
Table 6: Metrics for different search methods. We run search methods for 10 iters. each using gpt-3.5-turbo on HumanEval. Our method generates more diverse solutions and discovers the correct solution faster. We also compare against a genetic algorithm  (Romera-Paredes et al., 2024)
@@ 4.4 Solution diversity
We see in Table 6 that our proposed search method is able to propose more diverse candidate solutions with a lower semantic similarity score, while maintaining high quality search with by generating solutions with high validation scores. This shows that our methods help both exploration and exploitation without sacrificing too much of one for the other. Detailed stats shown in App. G.
@@ 4.5 Ablation on techniques
We performed an ablation study on the three introduced techniques as shown in Table 7 (and App. H), all of which enhanced performance and efficiency, with Scattering yielding the highest gains.
@@ 4.6 Verifier (Validation test) accuracy
Previous work emphasized the importance of a reliable verifier  (Liu et al., 2024a; Chen et al., 2022; Zhang et al., 2023a). In our case, we use self-generated validation tests noisy as a black-box verifer. Figure 13 shows the confusion matrix for whether self-generated validation tests accurately predict solution correctness. The false negative rate of 27.5% highlights misalignment between validation and ground truth tests. We performed an ablation study comparing performance when given different numbers of ground-truth tests as validation (Table 8). While finding the best verifier is not our focus, the results show a significant impact of verifier accuracy on pass@1. There are strong interactions between the search method and evaluation metric—just one ground-truth validation test dramatically improves performance in Table 5. However, the 33.75% inaccuracy of validation tests do not affect the pass@any rate much as shown in Table 8, which demonstrates the robustness of our method to validation noise. Detailed metrics shown in App. I.
@@ 4.7 Ablation on model
Fig. 14 shows weaker models scale better with our method, highlighting a trade-off: better-trained models scale worse with inference compute. This suggests inference-time scaling compensates for limited training quality, benefiting weaker models. Detailed stats are in App. F.
@@ 5 Additional Related work
Code generation with large language models. Some works focus on either training their own model  (Guo et al., 2024) or fine tuning existing models to adapt them towards code related tasks  (Roziere et al., 2023; Jain et al., 2023; Roziere et al., 2023). Recent literature has shown a flora of methods to improve LLM code generation performance during inference time, including agentic approaches  (Qian et al., 2023; Liu et al., 2024b), multi-agent approaches  (Tao et al., 2024; Hong et al., 2023; Islam et al., 2024), tool usage  (Zhang et al., 2024), and self-repair  (Le et al., 2023; Olausson et al., 2023a). There is a trade-off between scaling inference and training compute, with an optimal balance between them  (Wu et al., 2024; Sardana et al., 2024). Our work complements this by demonstrating the impact of enhanced solution diversity and exploration on scaling. While this work focuses on MCTS, our method can seamlessly integrate with advanced solution sampling methods during refinement  (Tang et al., 2024).
@@ 6 Conclusion
We have shown that framing code generation as an optimization task over the code space and applying Scattered Forest Search is highly effective. The simple yet powerful Scattering technique, which encourages the LLM to produce more diverse outputs, underscores the importance of exploration in optimization and search processes. This framework and these techniques can offer valuable insights for other code or language generation tasks.
The integration of search-based methods could significantly reduce computational costs while enhancing performance, making them attractive for large-scale deployments, especially in real-time applications or those requiring resource efficiency. This framework also lays the groundwork for future research into optimization strategies for language models, potentially leading to more advanced search algorithms, hybrid models, and novel techniques that push the limits of generative models.
@@ Appendix A Ethics Statement
Appendix B Problem setting and coding problem example
In the program synthesis task $\bm{x}=\langle\bm{p},\bm{H}\rangle$, the solver is provided with a problem prompt $\bm{p}$, described in natural language, which requests the generation of code to implement a specific functionality or object $\bm{s}$. The objective is to generate a solution $\bm{s}^{\prime}$ that passes all the hidden tests $\bm{H}$ designed to evaluate its correctness. These hidden tests $\bm{H}$ are not visible to the solver under any circumstances. In some cases, the solver may also have access to a set of validation tests $\bm{V}$, which they can use to verify their solution $\bm{s}^{\prime}$ before submitting it for evaluation against the hidden tests. Alternatively, the solver may generate their own validation tests $\bm{V}$ to test the functional correctness of $\bm{s}^{\prime}$. Both hidden and validation tests are typically structured as a set of assert statements validating specific functionalities of $\bm{s}$. A representative example of this problem setup is provided below.
@@ Appendix B Problem setting and coding problem example
A solution $\bm{s}^{\prime}$ is deemed correct if and only if it passes all hidden tests $\bm{H}$. Solvers are permitted up to $k$ attempts $[\bm{s}]_{k}$ to submit their solutions for evaluation. If at least one submission $\bm{s}^{*}$ satisfies all hidden tests, the task is considered solved. Given a set of tasks $\mathcal{X}$, the proportion of tasks $\langle\bm{p},\bm{H}\rangle\in\mathcal{X}$ successfully solved by the agent within $k$ submissions is referred to as the pass@k rate (Chen et al., 2021).
B.1 Justification for Problem Setup
In our problem setting, the solver is explicitly not allowed to access the hidden tests $\bm{H}$. This constraint ensures that the task closely mirrors real-world applications of program synthesis, where the solver must produce solutions based solely on the problem prompt $\bm{p}$. Such a setup aligns with platforms like LeetCode or conventional coding assessments, where test cases remain hidden to preserve the integrity of the evaluation process and discourage overfitting to known test sets.
This restriction is not only practical but also reflects the needs of end users in real-world scenarios. Writing comprehensive and correct unit tests is a time-consuming process, requiring substantial domain knowledge and effort. In practice, users would prefer tools that can autonomously generate correct implementations from a prompt without requiring additional test specifications or guidance. By requiring solvers to rely on self-generated validation tests or implicit reasoning to produce correct solutions, this setup encourages the development of tools capable of robust and generalized problem-solving.
By prioritizing correctness under these constraints, our framework evaluates an agent’s ability to understand the problem and synthesize accurate solutions in a realistic and user-centric manner, making it an ideal benchmark for program synthesis tasks.
Appendix C Methodology details
C.1 Traditional Monte Carlo Tree Search (MCTS)
Monte Carlo Tree Search (MCTS) is a widely used algorithm for decision-making in environments with large, complex state spaces, such as games and optimization problems. The algorithm incrementally builds a search tree by simulating possible outcomes and leveraging these simulations to balance exploration (visiting new states) and exploitation (refining knowledge about promising states). The MCTS process consists of four main stages: selection, expansion, simulation, and backpropagation.
C.1.1 Selection
The algorithm begins at the root node, corresponding to the current state $\bm{s}$, and iteratively selects child nodes using a selection policy designed to balance exploration and exploitation. The most commonly used selection policy is the Upper Confidence Bound for Trees (UCT), which evaluates each child state $\bm{s}^{\prime}$ and corresponding action (or direction) $\bm{d}$ according to the formula:
$UCT(\bm{s},\bm{d})=\widehat{Q}(\bm{s},\bm{d})+c\sqrt{\frac{\ln\left(\sum_{\bm{b}}n(\bm{s},\bm{b})\right)}{n(\bm{s},\bm{d})}},$ (6)
where $\widehat{Q}(\bm{s},\bm{d})$ is the estimated value of taking action $\bm{d}$ from state $\bm{s}$, $n(\bm{s},\bm{d})$ is the number of times this action has been visited, and $c$ is a hyperparameter controlling the exploration-exploitation trade-off. This equation encourages actions that either have high value (exploitation) or have been visited less frequently (exploration).
C.1.2 Expansion
After a leaf node is reached during selection, one or more child nodes representing potential future states $\bm{s}^{\prime}$ are added to the search tree. These states are typically generated by sampling actions or improvement directions $\bm{d}$ from a predefined set or policy. The added nodes represent unexplored areas of the search space.
C.1.3 Simulation
Once a new node is expanded, the algorithm performs a simulation or rollout from that state $\bm{s}^{\prime}$. In this step, a playout policy is used to sample a trajectory of states and actions until a terminal state or evaluation horizon is reached. The outcome of this simulation, denoted as $v(\bm{s}^{\prime})$, provides an estimate of the quality of the state $\bm{s}^{\prime}$. This value could represent a game outcome (e.g., win/loss), a performance metric (e.g., number of test cases passed), or another problem-specific objective.
C.1.4 Backpropagation
The results of the simulation are then propagated back up the search tree to update the value estimates $\widehat{Q}(\bm{s},\bm{d})$ for the nodes along the trajectory $\tau=[\bm{s}_{0},\bm{d}_{1},\bm{s}_{2},\dots,\bm{s}_{-1}]$. The value update is performed iteratively using a weighted average, which incorporates the new simulation result while retaining information from prior simulations:
$\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n})\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)}+\alpha_{n}\mathop{max}\{\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)},\widehat{Q}(\bm{s}_{i+1},\bm{d}_{i+2})^{(t+1)}\},$ (7)
where $\alpha_{n}$ is a weighting parameter that depends on the visit count $n(\bm{s},\bm{d})$, and $\mathop{max}$ ensures that only the best-performing solutions along the trajectory are prioritized.
C.2 Scattered forest search
Input: Language Model $\mathcal{L}$, Prompt $\bm{p}$, Number of iterations $N$, Branching factor $k$, Exploration parameter $c$
Output: Best solution $s^{*}$
Function Scattering($node$):
       Generate diverse directions $\{d_{1},d_{2},\dots,d_{k}\}$ using $\mathcal{L}$ conditioned on feedback for $node$;
       Save directions $\{d_{1},d_{2},\dots,d_{k}\}$ to $node.directions$, with $node.qvalue(d_{i})\leftarrow 0$;
       return $node.directions$;
Function Foresting($\mathcal{S}_{0}$):
       Initialize forest with random seed solutions $\{s_{1},s_{2},\dots,s_{n}\}$ sampled by $\mathcal{L}$;
       foreach $seed\in\mathcal{S}_{0}$ do
             Scattering($seed$); // Generate directions for each seed
       end foreach
      return forest $\mathcal{S}_{0}$;
Function Scouting($parent,direction,child,feedback,\mathcal{I}$):
       Update global insights $\mathcal{I}$ using feedback from applying $direction$ to $parent$, resulting in $child$;
Function SelectSeed($\mathcal{S}$):
       Select the most promising seed $s$ from $\mathcal{S}$ using UCT formula (Eq. 1);
       return $s$;
Function SelectDirection($node$):
       Select the most promising direction $d$ from $node.directions$ using UCT formula;
       return $d$;
Function Simulate($seed$):
       $parent\leftarrow seed,\tau\leftarrow\emptyset$;
       while $parent$ is not a leaf node do
             $direction\leftarrow$ SelectDirection($parent$);
             $child\leftarrow$ Child node corresponding to $direction$;
             $\tau.append((parent,direction,child))$;
             $parent\leftarrow child$;
       end while
      return leaf node and trajectory $\tau$;
Function Expand($parent,direction$):
       Generate new solution $\bm{s}^{\prime}$ using $\mathcal{L}$ conditioned on $\bm{s}$, $direction$, and feedback;
       Create new child node $child$ with solution $\bm{s}^{\prime}$ and feedback from validation tests on $parent$;
       return $child$;
Function Backpropagate($\tau,leaf,reward$):
       $leaf.qvalue[direction]\leftarrow reward$;
       foreach $(parent,direction,child)\in\tau$ (in reverse order) do
             Update $parent.qvalue[direction]$ using Eq. 2;
             $parent.visits[direction]\leftarrow parent.visits[direction]+1$;
       end foreach
begin
       Initialize global insights $\mathcal{I}\leftarrow\emptyset$;
       Initialize forest $\mathcal{S}_{0}\leftarrow$ Foresting();
       Generate validation tests $\bm{V}$ if not given;
       for $i\leftarrow 1$ to $N$ do
             $seed\leftarrow$ SelectSeed($\mathcal{S}_{0}$);
             $leaf,\tau\leftarrow$ Simulate($seed$);
             $direction\leftarrow$ SelectDirection($leaf$);
             $child\leftarrow$ Expand($leaf,direction$);
             $feedback,reward\leftarrow$ Execution feedback and validation score of $child$;
             Backpropagate($\tau,leaf,reward$);
             Scouting($leaf,direction,child,feedback,\mathcal{I}$);
             Scattering($child$); // Generate directions for the new node
       end for
      $s^{*}\leftarrow$ Best solution across all trees;
end
fAlgorithm 1 Scattered Forest Search (SFS) with Scouting
[…]
C.3 Algorithm Overview
SFS operates over multiple iterations, progressively refining the solution space using a search tree initialized with diverse seed solutions. For each iteration, SFS balances exploration and exploitation using the Upper Confidence Bound for Trees (UCT) formula to guide the search. The algorithm begins by initializing global insights $\mathcal{I}$, a forest of seed solutions $\mathcal{S}_{0}$, and validation tests $\bm{V}$. It then performs the following steps over $N$ iterations.
Initialization
  • Global Insights: A global memory $\mathcal{I}$ is initialized to store feedback and general insights gained during the search process, enabling knowledge sharing across branches.
  • Forest Creation: The algorithm generates a set of initial seed solutions $\mathcal{S}_{0}=\{\bm{s}_{1},\bm{s}_{2},\dots,\bm{s}_{n}\}$ by sampling diverse solutions using the language model $\mathcal{L}$. For each seed:
    • Direction Initialization: Diverse directions $\{d_{1},d_{2},\dots,d_{k}\}$ are generated for each seed solution by prompting $\mathcal{L}$ with feedback from the seed.
    • Q-Values: Q-values for each direction are initialized to 0, i.e., $node.qvalue(d_{i})\leftarrow 0$. These values are updated dynamically based on feedback during backpropagation.
Seed Selection
The algorithm selects the most promising seed solution $\bm{s}$ from the forest $\mathcal{S}$ using the UCT formula:
$UCT(\bm{s},\bm{d})=\widehat{Q}(\bm{s},\bm{d})+c\sqrt{\frac{\ln\left(\sum_{\bm{b}}n(\bm{s},\bm{b})\right)}{n(\bm{s},\bm{d})}},$ (8)
where $c$ is the exploration parameter, $n(\bm{s},\bm{d})$ is the visit count for direction (or seed) $\bm{d}$ at state $\bm{s}$, and $\widehat{Q}(\bm{s},\bm{d})$ is the estimated reward for that direction. This balances exploration of under-explored solutions and exploitation of high-reward solutions. Unvisited seeds/directions always have infinite UCT values, which means that we will always explore unvisited seeds or directions first.
Simulation
Starting from the selected seed $\bm{s}$, the algorithm simulates a trajectory $\tau=[(\bm{s}_{i},\bm{d}_{i},\bm{s}_{i+1})]$ by iteratively selecting the most promising direction $\bm{d}_{i}$ using the UCT formula. The simulation proceeds until a leaf node is reached.
Expansion
At the leaf node, the algorithm expands the search tree by generating a new child node:
$\bm{s}^{\prime}=\mathcal{L}(\bm{s},\bm{d},\text{feedback}),$ (9)
where $\mathcal{L}$ generates a new solution $\bm{s}^{\prime}$ by incorporating the parent solution $\bm{s}$, the improvement direction $\bm{d}$, and validation feedback. The newly generated solution $\bm{s}^{\prime}$ is validated, and the proportion of validation tests passed serves as the reward:
Reward=Number of tests passedTotal number of $\text{Reward}=\frac{\text{Number of tests passed}}{\text{Total number of tests}}.$ (10)
Backpropagation
The reward is backpropagated along the trajectory $\tau$:
$\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n})\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)}+\alpha_{n}\mathop{max}\{\widehat{Q}(\bm{s}_{i},\bm{d}_{i+1})^{(t)},\widehat{Q}(\bm{s}_{i+1},\bm{d}_{i+2})^{(t+1)}\},$ (11)
where $\alpha_{n}$ is a weighting parameter dependent on visit counts. The visit count $n(\bm{s},\bm{d})$ for each direction is incremented by 1 during backpropagation.
Scouting
The feedback $\bm{f}$ from applying a direction $\bm{d}$ to a parent node $\bm{s}$ is analyzed to extract general insights. These insights are stored in global memory $\mathcal{I}$ and reused to guide future direction generation. Scouting improves the search’s efficiency by leveraging knowledge gained from past iterations.
Scattering
For every expanded node, diverse new improvement directions $\{d_{1},d_{2},\dots,d_{k}\}$ are generated using $\mathcal{L}$. This encourages exploration of orthogonal regions of the solution space, reducing the likelihood of stagnation in local optima.
Final Selection
After $N$ iterations, the best solution $s^{*}$ across all trees in the forest is selected. The best solution is determined based on the highest reward, i.e., the proportion of validation tests passed.
C.4 Key Features of SFS
  • Q-Value Initialization: For every new direction, Q-values are initialized to 0 and dynamically updated during backpropagation.
  • Reward Definition: The reward is explicitly defined as the proportion of validation tests passed, providing a consistent measure of solution quality.
  • Diversity: By incorporating scattering and foresting, the algorithm ensures diversity in both initial seed solutions and branching directions, reducing the risk of stagnation in local optima.
  • Global Insights: The scouting mechanism enables the sharing of effective improvement strategies across the search space, enhancing the algorithm’s exploitation capabilities.
The SFS algorithm’s integration of scattering, foresting, and scouting represents a powerful approach to systematically navigate and optimize large solution spaces.
C.5 Algorithm complexity
The time complexity of the Scattered Forest Search (SFS) algorithm is primarily determined by the number of iterations $N$ and the traversal steps involved in each iteration.
Time Complexity:
At each iteration, the algorithm performs several steps: selecting seeds, simulating a trajectory, expanding nodes, and backpropagating rewards. In the worst-case scenario, the depth of the solution tree is proportional to $N$, resulting in $O(N)$ steps for traversal. Over $N$ iterations, this results in a total time complexity of $O(N^{2})$. Additionally:
  • Scouting and Scattering: The scouting and scattering steps at each iteration involve generating new directions and updating global insights, which have constant time complexity per iteration.
  • Comparison to Other Methods:
    • Line Search: Line search explores solutions sequentially without tree traversal, yielding a time complexity of $O(N)$.
    • Beam of Nodes (BoN): Similar to line search, BoN has a time complexity of $O(N)$ due to its focus on maintaining and updating a fixed number of candidate solutions.
    • Tree Search (MCTS): Traditional Monte Carlo Tree Search also has a time complexity of $O(N^{2})$, as it involves tree traversal similar to SFS.
Thus, the SFS algorithm shares the same $O(N^{2})$ time complexity as tree search but differs significantly in its use of global insights and scattering to improve solution diversity and efficiency.
Space Complexity:
  • SFS: The algorithm stores all generated solutions, resulting in a space complexity of $O(N)$. In addition, global insights $\mathcal{I}$ require constant space $O(1)$ since only the most recent updated insights need to be kept, and insights are kept within a fixed length.
  • Line Search: Line search has a space complexity of $O(1)$, as it retains only the most recent solution.
  • BoN and Tree Search: Both methods, like SFS, require $O(N)$ space to store all candidate solutions.
While the memory overhead for SFS is comparable to other tree-based methods, the efficient use of global insights enhances the algorithm’s performance without additional memory costs.
Practical Considerations:
Although SFS involves tree traversal, which contributes to the $O(N^{2})$ complexity, the computational bottleneck in practice is the solution generation process, typically dominated by LLM API calls. Thus, the tree traversal cost has a negligible impact on runtime in practical settings.
C.6 LLM request cost
Our algorithm requires two LLM-API calls (requests) per solution generated. One API call is used for code generation based on Scattering improvement directions. One API call is used for reflection after the code is generated and executed. The reflection step involves (1) updating the global insights based on the execution feedback (2) generating new Scattering directions based on the global insights and execution feedback. In practice, the two components of the reflection step can be separated into two API calls run in parallel to increase computational efficiency and code cleanness, which is how we implemented it.
Appendix D Scaling
Our method scales well with increased iterations. We show scaling curves for each dataset as shown in Figures 15, 16, 17. We also display scaling curves at higher budgets as shown in Figure 18. We see that performance mostly plateaus after the 20 solutions are generated. Moreover, other methods do not catch up to our method each with higher search budgets, suggesting how our method is able to fundamentally better explore the solution spaces, including solutions that other methods will not consider.